You have information which suggests the following equations for Demand and Supply: Qd = 200 -4P Qs = -4 + 6P What is the equilibrium price in this market? What is the equilibrium quantity in this market?
Answer #1

Answer:

Equilibrium quantity is the quantity where quantity demanded and quantity supplied are same. So that,

Qd=Qs

200 -4P=-4 + 6P

200+4=6P+4P

204=10P

So, P=20.4

Putting the value of P in the equation,

Qd = 200 -4P

Qd=200-4(20.4)

Qd=118.4

Equilibrium quantity=118.4